Three multiplied by the sum of 4 and a number is the same as 18 more than the number. Find the number.


If n is "the number," which equation could be used to solve for the number?

Answer by Theo(13342)

let x represent the number.

3 multiplied by the sum of 4 and the number is the same as 18 more than the number.

this statement translates to:

3 * (x + 4) = x + 18

solve for x as follows:

simplify the equation to get:
3x + 12 = x + 18
subtract 12 from both sides of the equation to get:
3x = x + 6
subtract x from both sides of the equation to get:
2x = 6
divide both sides of the equation by 2 to get:
x = 3

that should be your answer.

substitute the value of 3 for the number to see if this solves the original problem.

the original problem is:

Three multiplied by the sum of 4 and a number is the same as 18 more than the number. Find the number.

Assuming the number is 3, this statement then becomes:

Three multiplied by the sum of 4 and 3 is the same as 18 more than 3. Find the number.

18 more than 3 is equal to 21.

3 multiplied by the sum of 4 and 3 is 3 multiplied by 7 which is equal to 21.

This confirms the solution that the number is 3.